# Multi-Track MusicLDM: Towards Versatile Music Generation with Latent Diffusion Model

### v1

@@ 2 Background
Arrangement generation. The most similar work to ours is the commercial product JEN-1 Composer [45], a framework that uses a latent diffusion-based music generation system for versatile multi-track music generation. Different from our approach, they use a latent space obtained directly from audio with their own pre-trained autoencoder model for audio reconstruction, following the Jen-1 architecture proposed in [21]. Additionally, the dataset used is also different from ours, as they used their own private dataset, which is larger and of higher quality than the publicly available Slakh2100 [27] dataset that we used. Unfortunately, as it is a commercial product, they do not share code, datasets, or provide a free API for audio generation, making direct comparison to our work impossible. Other recently published works have explored music-to-music and arrangement-like music generation. STEMGEN [32] presented an alternative paradigm for music generation by introducing a model that can listen and respond to musical context. Different from ours, STEMGEN uses a transformer language model-like architecture on hierarchical discrete representations from VQ-VAEs [41] to model musical tracks. Some other works approach the arrangement task with a one-to-many or many-to-one paradigm. For example, [33] generates bass accompaniment using audio autoencoders. On the other hand, SingSong [4], utilizing an architecture similar to AudioLM [1], proposed a method for generating instrumental music to accompany input vocals and demonstrated promising results. However, these works primarily concentrate on one instrument, offering limited versatility. Another work that serves as our baseline, MSDM [28], introduced a novel diffusion-based, multi-source generative framework trained via denoising score-matching [39, 40]. MSDM is capable of synthesizing music, creating arrangements, and separating musical sources while operating in the raw waveform domain. While MSDM is notable for its flexibility, it exhibits limitations in audio quality and musical coherence. Interestingly, although our model surpasses MSDM in most tasks, it struggles with source separation—a limitation we attribute to its operation within a latent space where mixtures and stems do not maintain a linear relationship. This insight is prompting our future research in this direction.
@@ 3.2 Conditional Generation
Additionally, to enable more controllable generation, LDM models often employ classifier-free guidance (CFG) [10]. CFG is a technique in diffusion models that enhances control over the adherence to conditioning information during inference. This is achieved by randomly dropping the conditioning information during training, thereby simultaneously training both conditional and unconditional versions of the LDM model. In the inference time, the strength of the conditioning can be modulated by the CFG weight $\hat{\epsilon}=w\epsilon_{u}+(1-w)\epsilon_{c}$, where $w$ is the guidance scale weight that balances the model’s unconditional $\epsilon_{u}$ and conditional $\epsilon_{c}$ predictions.

### v2

@@ 2 Related Work
Arrangement generation. A commercial product JEN-1 Composer [47] introduced generating musical arrangements in audio domain using a latent diffusion-based music generation system. However, unlike our approach, JEN-1 Composer is not multi-track and uses a latent space obtained directly from audio with its own pre-trained autoencoder model for audio reconstruction, following its predecessor, the Jen-1 architecture proposed in [21]. Additionally, the dataset used is also different from ours, as they used their own private dataset, which is larger and of higher quality than the publicly available Slakh2100 [27] dataset that we used. Unfortunately, as it is a commercial product, they do not share code, datasets, or provide a free API for audio generation, making direct comparison to our work impossible. Other recently published works have explored music-to-music and arrangement-like music generation. STEMGEN [33] presented an alternative paradigm for music generation by introducing a model that can listen and respond to musical context. Different from ours, STEMGEN uses a transformer language model-like architecture on hierarchical discrete representations from VQ-VAEs [42] to model musical tracks. Diff-A-Riff [30] presented a non-multi track approach to generating instrument accompaniments using diffusion models and incorporates control mechanisms like both text-driven and audio-driven CLAP conditioning. The most similar work to ours is [45], which was published during the review process of our paper. This study also explored multi-track music generation using the same Slakh2100 dataset and reported results that are very similar to ours.
Some other works approach the arrangement task with a one-to-many or many-to-one paradigm. For example, [34] generates bass accompaniment using audio autoencoders. On the other hand, SingSong [4], utilizing an architecture similar to AudioLM [1], proposed a method for generating instrumental music to accompany input vocals and demonstrated promising results. However, these works primarily concentrate on one instrument, offering limited versatility. Another work that serves as our baseline, MSDM [28], introduced a novel diffusion-based, multi-source generative framework trained via denoising score-matching [40, 41]. MSDM is capable of synthesizing music, creating arrangements, and separating musical sources while operating in the raw waveform domain. While MSDM is notable for its flexibility, it exhibits limitations in audio quality and musical coherence. Interestingly, although our model surpasses MSDM in most tasks, it struggles with source separation—a limitation we attribute to its operation within a latent space where mixtures and stems do not maintain a linear relationship. This insight is prompting our future research in this direction.
@@ 3.2 Conditional Generation
Additionally, to enable more controllable generation, LDM models often employ classifier-free guidance (CFG) [10]. CFG is a technique in diffusion models that enhances control over the adherence to conditioning information during inference. This is achieved by randomly dropping the conditioning information during training, thereby simultaneously training both conditional and unconditional versions of the LDM model. In the inference time, the strength of the conditioning can be modulated by the CFG weight $\hat{\epsilon}=w\epsilon_{c}+(1-w)\epsilon_{u}$, where $w$ is the guidance scale weight that balances the model’s unconditional $\epsilon_{u}$ and conditional $\epsilon_{c}$ predictions.
@@ 6 Conclusion and Future Work
In the future, we also see potential for MT-MusicLDM to become an interactive tool for both amateur musicians and educational purposes. For example, it could be developed as a web-based system where users can generate full musical pieces or arrangements based on selected tracks or styles, helping them explore different musical directions. The model could also offer the ability to generate similar tracks or continue existing ones, making it an intuitive tool for creativity and learning.